# RLAIF-V: Aligning MLLMs through Open-Source AI Feedback for Super GPT-4V Trustworthiness

### v1

@@ 3.2 Main Results
The main experimental results are reported in Table 1, from which we observe that: (1) RLAIF-V achieves state-of-the-art performance in trustworthiness among open-source models and even surpasses proprietary models such as GPT-4V. The framework significantly reduces the object hallucination rate of LLaVA 1.5 and OmniLMM by 82.9% and 71.0% relative points on Object HalBench. For the overall hallucination rate, RLAIF-V 12B achieves 29.5% on MHumanEval, surpassing GPT-4V by a large margin. The reduction of hallucination is consistent among multiple benchmarks including MMHal-Bench, AMBER, and RefoMB. (2) RLAIF-V achieves promising performance in response helpfulness, where the results on LLaVA Bench and MMStar are improved compared to the base model. This shows that RLAIF-V can enhance the trustworthiness of MLLMs without sacrificing the performance of other tasks. (3) Using OmniLMM as both the instruction model and the labeler model, RLAIF-V 12B achieves significant hallucination reduction on multiple benchmarks and comparable helpfulness. Remarkably, RLAIF-V 12B outperforms GPT-4V in trustworthiness on Object HalBench, MHumanEval, AMBER, and RefoMB, by substantial margins. The results demonstrate a promising path to achieve self-alignment of leading-edge MLLMs.
Data ObjHal. AMBER
Resp. $\downarrow$ Ment. $\downarrow$ Acc. F1
RLHF-V [58] 28.5 12.3 76.4 84.6
Adapted 25.7 11.8 73.3 83.0
Ours 10.1 4.7 80.1 86.1
@@ 3.3 Analysis
We conduct analysis on the framework considering the following research questions: (1) How effective is the deconfounded candidate response generation strategy for improving feedback quality? (2) Can we directly acquire satisfactory feedback from open-source MLLMs without the divide-and-conquer approach? (3) What is the advantage of learning from feedback in an iterative manner? (4) Can RLAIF-V be used with other sources of feedback together? (5) Can feedback data collected for one model with RLAIF-V be adopted to enhance the trustworthiness of other MLLMs?
Deconfounded strategy enables better learning efficiency. To quantify the advantage of the deconfounded candidate response generation strategy, we conduct an experiment based on the RLHF-V dataset [58]. We compare the performance of model trained under three different settings: (1) RLHF-V, the model is directly aligned with fine-grained correctional human feedback data; (2) Ours, we collect high-quality feedback from LLaVA-NeXT based on original multimodal instructions using the RLAIF-V framework; (3) Adapted, we replace the preferred responses generated under the deconfounded strategy with original human annotations.
From experimental results in Table 3.2, we observe that model trained with our deconfounded responses achieves the best performance on both long-form description and discriminative tasks. Changing the preferred response with high quality human annotated response, though improving the feedback precision and response quality, exhibits significant performance loss. We hypothesize this action introduces more non-robust shallow patterns into the training data and thus harms the learning efficiency. Moreover, performance of our method even surpasses training on fine-grained human-annotated correctional feedback by a large margin.
Table 3: Performance comparison of different feedback collection methods. We conduct the experiment on various labeler models. ObjHal.: Object HalBench.
Method Labeler Labeler Size ObjHal. AMBER
Resp. $\downarrow$ Ment. $\downarrow$ Acc. F1
Self Rewarding [61] LLaVA 1.5 7B 48.4 24.6 54.5 48.3
OmniLMM 12B 46.7 23.9 57.5 53.2
LLaVA-NeXT 34B 58.4 28.3 74.5 78.5
Ours LLaVA 1.5 7B 39.5 19.6 77.4 81.3
OmniLMM 12B 33.0 17.5 75.2 78.2
LLaVA-NeXT 34B 20.6 10.4 80.5 86.0
Divide-and-conquer enhances feedback quality. We compare our divide-and-conquer strategy with direct self-rewarding [61] by replacing only the implementation of response evaluation process. Specifically, self-rewarding asks the labeler model to generate an overall quality score of each candidate response with a long prompt introducing multiple criteria. To bolster the reliability, we adopt three labeler models from weak to strong and present the results in Table 3. We observe that simply asking open-source models to generate an overall assessment of responses yields unsatisfactory results and can even harm the performance. In contrast, our method not only improves performance but also demonstrates increasing efficacy as the size of the labeler model grows, showing promising scalability and compatibility.
Figure 3: Comparison of iterative and non-iterative alignment. We report overall win rate compared with GPT-4V on RefoMB and response-level non-hallucination rate on Object HalBench.
Iterative alignment mitigates distribution shift problem. To quantify the effectiveness of iterative alignment, we compared the results of our iterative alignment and non-iterative naive direct preference optimization in Figure 3. Considering the scale of data, we conduct this experiment under the RLAIF-V 12B setting to save cost with a smaller labeler model. We observe that the performance of the non-iterative method quickly saturates and the effectiveness of data cannot be fully utilized, while iterative alignment continues to reduce object hallucination and improve overall win rate compared with GPT-4V on RefoMB.
Figure 4: (a) Results of combining RLAIF-V with other sources of feedback. (b) Hallucination reduction of other MLLMs with data from the first training iteration of RLAIF-V 12B (i.e., using OmniLMM as instruction model). We report the response-level hallucination rate reduction on Object HalBench for generative hallucination and error rate reduction on AMBER for discriminative hallucination.
@@
RLAIF-V is complementary with existing feedback learning methods. Besides collecting feedback using models as labelers, multiple existing works generate feedback based on heuristic rules or human annotation. We explore the possibility of combining RLAIF-V with other sources of feedback. Results in Figure 4 show that heuristically collected feedback from HA-DPO [67] and human annotated feedback from RLHF-V can further improve the trustworthiness, indicating RLAIF-V is complementary with other types of feedback.
@@
RLAIF-V produces generalizable high-quality feedback. We train different models with the feedback collected during the first iteration of training RLAIF-V 12B. Specifically, we train LLaVA 1.5 7B [31], LLaVA 1.5 13B [31], MiniCPM-V [43] and MiniCPM-V 2 [43] with direct preference optimization and report the trustworthiness improvement in Figure 4 (b). We observe that data collected from OmniLMM (as both instruction model and labeler model) with RLAIF-V can effectively reduce the hallucination of other MLLMs on different benchmarks. Notably, the improvement can be even more significant compared with the OmniLMM which generates the candidate responses. The results demonstrate that feedback data collected with RLAIF-V are generalizable to improve trustworthiness of different MLLMs.
3.4 Case Study
To provide an intuitive understanding and comparison of different models, we provide qualitative results in Figure 5. We show cases of two RLAIF-V models: (1) In the first example, both RLAIF-V 7B and GPT-4V correctly identify the relationships between characters in the image as colleagues or members with a common purpose. However, GPT-4V incorrectly claims that the person talking to the rest of the group is standing. (2) In the second example, RLAIF-V 12B exhibits superior trustworthiness level compared with GPT-4V, without introducing any hallucination. On the other hand, GPT-4V incorrectly mentions the blue outline and darker blue URLs which do not exist in the image. We refer readers to the Appendix for more qualitative results.
@@ 4 Conclusion
Aligning models with human preference is a critical target. In this work, we present RLAIF-V, a novel framework that enhances the trustworthiness of MLLMs through open-source AI feedback. Comprehensive experimental results show that our models achieve state-of-the-art performance in both generative and discriminative trustworthiness. In this work, we propose a deconfounded sampling and divide-and-conquer strategy to improve the efficiency and quality of feedback. By iteratively aligning the model with such feedback, the trustworthiness can be substantially improved without sacrificing performance on other tasks. In the future, we will explore collecting more complex feedback from models to improve logical reasoning and complex task solving capabilities.
Appendix A Related Work
@@
Feedback Collection for MLLMs. Feedback quality is one of the most important factors for models to align with human preferences. Early works mainly collect high-quality feedback through human labelers which is costly and limited compared with the widespread misalignment problem [58, 49]. To this end, collecting feedback from AI serves an alternative to get rid of human intervention and provides a promising way to guide super-intelligent models beyond human performance [9]. However, existing methods simply distill feedback for MLLMs from proprietary models like GPT-4V, which rely on the superiority of proprietary model over the student model which uses the feedback to improve itself [26]. The concurrent HSA-DPO [56] asks GPT-4 [41] and GPT-4V [42] to detect hallucination from 6k image descriptions and use the output to train a 40B detector model for hallucination detection. It then applies a 34B re-writer model to re-write hallucinated sentences to form preference pairs. FGAIF [20] asks ChatGPT to split the response into sub-sentences and classify them into either object-existence or attribute or relation relevant which are further used to collect feedback from the LLaVA 1.5 13B to get a score of each response. These approaches still depend on strong proprietary models and only tackle MLLM hallucination on the image captioning task regarding three kinds of object-related hallucination. RLAIF-V, on the other hand, strengthens MLLMs with feedback on a diverse range of tasks (e.g., visual question answering [38], scene text understanding [48] and image captioning [30]) under a fully open-source setting. HA-DPO [67], POVID [69] and IBD [71] heuristically construct comparison pairs by either distorting the image or editing the model response. FDPO [16] employs human annotators to collect span-level fine-grained feedback to reduce the hallucination of MLLMs.
@@ B.1 GPT-4 as Evaluator
Table 4: The number of instructions in each category of RefoMB.
@@ B.2.1 Image and Description Collection
Annotating a comprehensive description avoiding missing any important content in the image in a single turn can be overly challenging for even experienced annotators. In order to achieve reliable high content coverage and accuracy, we devise a three-step process as follows: (1) We first employ GPT-4V to generate detailed descriptions based on six different prompts listed in Table 8, where these prompts focus on different aspects of the same image. (2) We then merge these image descriptions, each with its own focus, to generate a draft comprehensive description by asking GPT-4 with prompt in Table 8. (2) Such a draft can be limited in both coverage and accuracy, so we ask human annotators to add more details and correct errors for each draft description. (3) To further ensure the comprehensiveness of image descriptions, each description is verified by at least two graduate students. (4) To ensure the accuracy and completeness of the annotations, the descriptions underwent a minimum of three rounds of additions and modifications. Specifically, the annotation price is 10 dollars per hour in average.
@@ B.3.1 Reliability Analysis of RefoMB
To explore the reliability of our evaluation results, we conduct an experiment comparing the human agreement of RefoMB with widely used MMHal-Bench [49], which all utilize GPT-4 as the evaluator. Specifically, we use these benchmarks to assess the performance of six commonly employed MLLMs, including LLaVA 1.5 [31], LLaVA-NeXT [32], GPT-4V [42], OmniLMM [43], RLAIF-7B and RLAIF-12B. For each instruction in every benchmark, we collect $302\times\binom{6}{2}=$ 30 response pairs by combining outputs generated by different models. Then, we uniformly sample 100 pairs from each benchmark and collect corresponding win-lose-tie decisions. For MMHal-Bench, which assigns absolute scores for each response, we compare the score value of two responses to get the decision. We then ask the human annotator to classify each evaluation result into “agree” or “disagree” and present the results in Table B.3.1. We observe that RefoMB exhibits both higher reliability and more evenly distributed win and lose counts.
@@ B.4 Example of Evaluation Results
To provide a more intuitive understanding of evaluation results on different tasks. We present an example of evaluation results from RefoMB in Figure 11.
@@ B.6 RefoMB Test Split Evaluation Results
LLaVA 1.5 [31] 56.2 20.8 28.9 13.2 12.5 4.2 13.6 4.5 21.4 14.3 33.3 4.2 0.0 0.0 37.5 25.0 30.8 12.1
LLaVA 1.5 + RLAIF-V 70.8 22.9 42.1 23.7 58.3 45.8 22.7 18.2 39.3 14.3 33.3 8.3 50.0 0.0 37.5 0.0 47.5 20.7
OmniLMM [43] 64.6 47.9 63.2 13.2 54.2 33.3 18.2 18.2 42.9 7.1 33.3 20.8 0.0 0.0 37.5 0.0 48.0 23.7
OmniLMM + RLAIF-V 81.2 45.8 68.4 44.7 75.0 25.0 22.7 9.1 64.3 28.6 50.0 29.2 0.0 0.0 62.5 37.5 62.1 32.8
@@
LLaVA 1.5 [31] 54.1 15.6 36.0 17.0 27.3 10.6 32.1 16.1 34.2 25.0 42.2 18.8 0.0 0.0 36.4 0.0 38.1 15.9
+ RLAIF-V 59.8 25.4 49.0 14.0 42.4 18.2 32.1 16.1 34.2 25.0 40.6 26.6 31.2 6.2 45.5 22.7 45.0 20.7
OmniLMM [43] 51.6 20.5 54.0 22.0 45.5 9.1 35.7 25.0 53.9 23.7 46.9 17.2 18.8 0.0 54.5 9.1 48.5 18.8
+ RLAIF-V 72.1 31.1 59.0 36.0 51.5 22.7 41.1 23.2 53.9 30.3 56.2 35.9 31.2 18.8 63.6 18.2 57.5 29.7

### v2

@@ 3.2 Main Results
The main experimental results are reported in Table 1, from which we observe that: (1) RLAIF-V achieves state-of-the-art performance in trustworthiness among open-source models and even surpasses proprietary models such as GPT-4V. The framework significantly reduces the object hallucination rate of LLaVA 1.5 and OmniLMM by 80.7% and 76.8% relative points on Object HalBench. For the overall hallucination rate, RLAIF-V 12B achieves 35.6% on MHumanEval, surpassing GPT-4V by a large margin. The reduction of hallucination is consistent among multiple benchmarks including MMHal-Bench, AMBER, and RefoMB. (2) RLAIF-V achieves promising performance in response helpfulness, where the results on MMStar are improved compared to the base model. This shows that RLAIF-V can enhance the trustworthiness of MLLMs without sacrificing the performance of other tasks. (3) Using OmniLMM as both the instruction model and the labeler model, RLAIF-V 12B achieves significant hallucination reduction on multiple benchmarks and comparable helpfulness. Remarkably, RLAIF-V 12B outperforms GPT-4V in trustworthiness on Object HalBench, MHumanEval, AMBER, and RefoMB, by substantial margins. The results demonstrate a promising path to achieve self-alignment of leading-edge MLLMs. (4) Self-feedback guidance improves the trustworthiness of both RLAIF-V 7B and RLAIF-V 12B on multiple benchmarks with best-of-N selection, demonstrating the effectiveness of RLAIF-V reward at inference-time.
3.3 Ablation Study
To investigate the contribution of different components in RLAIF-V, we perform an ablation study.
Ablation of response generation approach. To quantify the advantage of the deconfounded candidate response generation strategy, we conduct an experiment based on the RLHF-V dataset [66]. We compare the performance of model trained under three settings: (1) RLHF-V, the model is directly aligned with human feedback data; (2) RLAIF-V, we collect high-quality feedback from LLaVA-NeXT based on original multimodal instructions in RLHF-V dataset using the RLAIF-V framework; (3) RLAIF-V w/o deconfounding, we replace the preferred responses generated under the deconfounded strategy with original human annotations.
From experimental results in Table 2, we observe that the model trained with our deconfounded responses achieves the best performance on both tasks. Changing preferred responses with high-quality human annotated responses, though improving the feedback precision and response quality, exhibits significant performance loss. We hypothesize this action introduces more non-robust shallow patterns into the training data and thus harms the learning efficiency. Moreover, the performance of our method even surpasses training on human-annotated correctional feedback by a large margin. After analyzing the composition of the RLHF-V dataset, we find it only includes a limited selection of models [66] which share limited hallucination distribution similarity with LLaVA 1.5 7B. As a result, the effectiveness of the dataset is significantly diminished. We argue this phenomenon further enhances the importance of the RLAIF-V framework which can efficiently generate high-quality feedback data for any MLLM. We list more details about the hallucination distribution and RLHF-V dataset composition in the Appendix D.
Effect of divide-and-conquer strategy. We compare our divide-and-conquer strategy with direct self-rewarding [69] by replacing only the implementation of response evaluation process. Specifically, self-rewarding asks the labeler model to generate an overall quality score of each candidate response with a long prompt introducing multiple criteria. We assess the human agreement of generated response pairs by the ratio of $(y_{w},y_{l})$ which evaluators agree that $y_{w}\geqslant y_{l}$. Based on results in Table 3, we observe that simply asking open-source models to generate an overall assessment of responses yields unsatisfactory results due to poor feedback quality. In contrast, our method with the divide-and-conquer strategy significantly improves the feedback quality and overall performance on both discriminative and generative tasks. Moreover, we also compare RLAIF-V feedback data with VL-Feedback [28] which collects high-quality feedback from GPT-4V. Results show RLAIF-V achieves higher data quality with a novel divide-and-conquer strategy and better performance by training on the same amount of data.
Figure 3: Results of combining RLAIF-V with other feedback. We report the response-level no-hallucination rate on Object HalBench for generative trustworthiness.
Figure 4: Hallucination reduction of other MLLMs with data from the first training iteration of RLAIF-V 12B. We report the response-level hallucination rate reduction on Object HalBench for generative hallucination and AMBER error rate reduction for discriminative hallucination.
Figure 5: Inference-time scaling across different models. We report the response level no-hallucination rate on Object HalBench for generative trustworthiness. Reference PPL: using perplexity from OmniLMM.
@@ 3.4 Analysis
We conduct analysis on the framework considering the following research questions: (1) Is RLAIF-V compatible with other sources of feedback together? (2) Can feedback data collected for one model with RLAIF-V be adopted to enhance the trustworthiness of other MLLMs? (3) How RLAIF-V reward work for inference-time scaling?
@@
RLAIF-V is complementary with existing feedback collection methods. Besides collecting feedback using models as labelers, multiple existing works generate feedback based on heuristic rules or human annotation. We explore the possibility of combining RLAIF-V with other sources of feedback. Results in Figure 3 show that heuristically collected feedback from HA-DPO [76] and human annotated feedback from RLHF-V can further improve the trustworthiness, indicating RLAIF-V is complementary with other types of feedback.
@@
RLAIF-V produces generalizable high-quality feedback. We train different models with the feedback collected during the first iteration of training RLAIF-V 12B. Specifically, we train LLaVA 1.5 7B [33], LLaVA 1.5 13B [33], MiniCPM-V [46], MiniCPM-V 2 [46]. with direct preference optimization and report the trustworthiness improvement in Figure 4. We observe that data collected from OmniLMM (as both the instruction model and labeler model) with RLAIF-V framework can effectively reduce the hallucination of other MLLMs on different benchmarks. Notably, the improvement can be even more significant compared with the OmniLMM which generates the candidate responses. The results demonstrate that feedback from RLAIF-V is generalizable to improve the trustworthiness of different MLLMs.
RLAIF-V reward continuously improves MLLM trustworthiness at inference-time. We explore the effectiveness of RLAIF-V 12B reward on different open-source MLLMs. As results shown in Figure 5, RLAIF-V reward consistently improves the generative trustworthiness of both LLaVA 1.5 7B and Qwen-VL-Chat. We compare the improvement with two baselines which use the perplexity (PPL) of RLAIF-V 12B model or OmniLMM and observe that RLAIF-V reward achieves significantly better results. We also analyze the average length of best-of-N selected responses compared with naive sampled responses and demonstrate that the simple length-normalization method effectively tackles the bias of preferring shorter responses which might cause significant information loss. Specifically, the average length difference count by words is increased from -7.7 (w/o length-normalization) to +3.9 when using RLAIF-V 12B reward for best-of-64 setting of LLaVA 1.5 7B.
4 Related Works
We introduce the most related background works in this section and refer readers to the Appendix for a more detail review of related works.
Learning from Feedback.
Learning from feedback is one of the core techniques in developing advanced LLMs [9, 68, 57] and MLLMs [66, 56, 28, 78], which aligns the model with human preference. Proximal policy optimization (PPO) [52] is recognized as the major technique to directly align models with human preferences through training a reward model on pairwise comparisons of model responses. Rafael et al. [48] propose direct preference optimization to stabilize the training of PPO and is widely adopted by the community recently. However, most multimodal feedback learning methods only utilize the simplicity and training stability of DPO while omits the important fact that DPO actually trains an optimal reward model. As a result, without explore the effectiveness of the continuous rewards, these methods gets suboptimal outcomes.
Feedback Collection for MLLMs. Feedback quality is one of the most important factors for models to align with human preferences. Early works mainly collect high-quality feedback through human labelers which is costly and limited compared with the widespread misalignment problem [66, 56, 16]. To this end, collecting feedback from AI serves an alternative to get rid of human intervention and provides a promising way to guide super-intelligent models [9]. However, existing methods simply distill feedback for MLLMs from proprietary models like GPT-4V, which rely on the superiority of proprietary model over the student model which uses the feedback to improve itself [28]. The concurrent HSA-DPO [63] asks GPT-4 [44] and GPT-4V [43] to detect hallucination from 6k image descriptions. FGAIF [22] asks ChatGPT to split the response into sub-sentences and classify them into either object-existence or attribute or relation relevant. These approaches still depend on strong proprietary models and only tackle MLLM hallucination on the image captioning task regarding three kinds of object-related hallucination. RLAIF-V, on the other hand, strengthens MLLMs with feedback on a diverse range of tasks (e.g., visual question answering [40], scene text understanding [53] and image captioning [32]) under a fully open-source setting. HA-DPO [76], POVID [78], AMP [74] and BPO [47] heuristically construct comparison pairs by either distorting the image, editing the model response or pairing models with different performance.
[…]
@@ 5 Conclusion
Aligning models with human preference to reduce MLLM hallucination is a critical target. In this work, we present RLAIF-V, a novel framework that enhances the trustworthiness of MLLMs through open-source AI feedback. Comprehensive experimental results show that our models achieve state-of-the-art performance in both generative and discriminative trustworthiness. We propose a deconfounded sampling and divide-and-conquer strategy to improve the efficiency and quality of feedback. By aligning the model with such high-quality feedback, the trustworthiness can be substantially improved without sacrificing performance on other tasks. Moreover, we propose novel self-feedback guidance for inference-time scaling using the aligned model itself and a simple length-normalization strategy to tackle the bias towards shorter responses. We also demonstrate that feedback generated via the RLAIF-V framework is generalizable to different MLLMs. In the future, we will explore collecting more complex feedback from models to improve logical reasoning and complex task-solving capabilities.
Supplementary Material
Appendix A Extended Related Work
@@
Feedback Collection for MLLMs. Feedback quality is one of the most important factors for models to align with human preferences. Early works mainly collect high-quality feedback through human labelers which is costly and limited compared with the widespread misalignment problem [66, 56]. To this end, collecting feedback from AI serves an alternative to get rid of human intervention and provides a promising way to guide super-intelligent models beyond human performance [9]. However, existing methods simply distill feedback for MLLMs from proprietary models like GPT-4V, which rely on the superiority of proprietary model over the student model which uses the feedback to improve itself [28]. The concurrent HSA-DPO [63] asks GPT-4 [44] and GPT-4V [43] to detect hallucination from 6k image descriptions and use the output to train a 40B detector model for hallucination detection. It then applies a 34B re-writer model to re-write hallucinated sentences to form preference pairs. FGAIF [22] asks ChatGPT to split the response into sub-sentences and classify them into either object-existence or attribute or relation relevant which are further used to collect feedback from the LLaVA 1.5 13B to get a score of each response. These approaches still depend on strong proprietary models and only tackle MLLM hallucination on the image captioning task regarding three kinds of object-related hallucination. RLAIF-V, on the other hand, strengthens MLLMs with feedback on a diverse range of tasks (e.g., visual question answering [40], scene text understanding [53] and image captioning [32]) under a fully open-source setting. HA-DPO [76], POVID [78] and BPO [47] heuristically construct comparison pairs by either distorting the image or editing the model response. FDPO [16] employs human annotators to collect span-level fine-grained feedback to reduce the hallucination of MLLMs.
@@ B.1 GPT-4 as Evaluator
Total 85 69 45 39 52 44 11 15 360
Table 4: The number of instructions in each category of RefoMB.
@@ B.2.1 Image and Description Collection
Annotating a comprehensive description avoiding missing any important content in the image in a single turn can be overly challenging for even experienced annotators. In order to achieve reliable high content coverage and accuracy, we devise a three-step process as follows: (1) We first employ GPT-4V to generate detailed descriptions based on six different prompts listed in Table 8, where these prompts focus on different aspects of the same image. (2) We then merge these image descriptions, each with its own focus, to generate a draft comprehensive description by asking GPT-4 with prompt in Table 8. (2) Such a draft can be limited in both coverage and accuracy, so we ask human annotators to add more details and correct errors for each draft description. (3) To further ensure the comprehensiveness of image descriptions, each description is verified by at least two graduate students. (4) To ensure the accuracy and completeness of the annotations, the descriptions underwent a minimum of three rounds of additions and modifications. Specifically, the annotation price is 10 dollars per hour in average and only annotators with an English proficiency equivalent to a TOEFL score of 110 or higher are involved.
@@ B.3.1 Reliability Analysis of RefoMB
To explore the reliability of our evaluation results, we conduct an experiment comparing the human agreement of RefoMB with widely used MMHal-Bench [56], which all utilize GPT-4 as the evaluator. Specifically, we use these benchmarks to assess the performance of six commonly employed MLLMs, including LLaVA 1.5 [33], LLaVA-NeXT [35], GPT-4V [43], OmniLMM [46], RLAIF-7B and RLAIF-12B. For each instruction in every benchmark, we collect $302\times\binom{6}{2}=$ 30 response pairs by combining outputs generated by different models. Then, we uniformly sample 100 pairs from each benchmark and collect corresponding win-lose-tie decisions. For MMHal-Bench, which assigns absolute scores for each response, we compare the score value of two responses to get the decision. We then ask the human annotator to classify each evaluation result into “agree” or “disagree” and present the results in Table 5. We observe that RefoMB exhibits both higher reliability and more evenly distributed win and lose counts.
@@ B.4 Example of Evaluation Results
To provide a more intuitive understanding of evaluation results on different tasks. As shown in Figure 11, we show a case of evaluation result from RefoMB.
@@ B.6 RefoMB Test Split Evaluation Results
LLaVA 1.5 [33] 54.9 20.1 40.4 18.4 34.7 23.6 15.2 4.6 33.3 13.1 29.2 11.1 0.0 0.0 41.7 29.2 36.9 16.2
+ RLAIF-V 68.2 27.6 51.3 30.9 51.0 30.2 26.1 18.2 42.0 19.6 38.5 11.5 16.7 0.0 15.6 0.0 47.2 22.5
+ RLAIF-V BoN 70.3 28.6 57.2 26.3 65.6 40.6 36.4 22.7 63.4 25.9 41.7 10.4 12.5 0.0 31.3 0.0 55.7 24.4
OmniLMM [46] 55.6 27.4 50.9 14.9 56.2 22.2 26.5 19.7 33.3 16.1 40.3 15.3 11.1 0.0 43.8 0.0 44.7 18.5
+ RLAIF-V 75.0 45.8 57.9 29.0 66.7 41.7 31.8 4.6 57.1 17.9 45.8 29.2 33.3 0.0 62.5 0.0 58.1 28.3
+ RLAIF-V BoN 84.2 34.6 62.6 42.6 75.8 34.2 35.5 10.0 57.1 25.0 50.8 33.3 20.0 0.0 62.5 22.5 62.9 30.3
@@
LLaVA 1.5 [33] 50.0 15.6 31.0 18.0 22.7 6.1 33.9 19.6 36.8 22.4 42.2 15.6 12.5 0.0 40.9 9.1 36.8 15.5
+ RLAIF-V 59.8 18.0 46.0 21.0 39.4 12.1 37.5 17.9 39.5 29.0 35.9 15.6 31.3 0.0 36.4 9.1 44.4 18.2
+ RLAIF-V BoN 66.4 20.5 51.0 25.0 47.0 12.1 35.7 16.1 38.2 25.0 37.5 18.8 37.5 0.0 54.6 0.0 48.7 18.8
OmniLMM [46] 54.1 15.6 56.0 25.0 43.9 6.1 33.9 14.3 35.5 25.0 48.4 17.2 6.3 0.0 36.4 0.0 45.4 16.5
+ RLAIF-V 65.6 26.5 55.0 29.7 54.0 18.7 32.1 16.7 56.6 39.5 55.7 25.0 29.2 6.3 63.6 21.2 54.8 25.9
+ RLAIF-V BoN 65.8 32.2 61.3 31.0 53.5 14.1 40.5 15.5 56.6 31.1 53.6 22.4 27.1 6.3 71.2 18.2 56.9 25.2